### v1

# Adaptive refinement for eigenvalue problems based on an associated source problem

Adaptive refinement for eigenvalue problems based on an associated source problem
Stefano Giani, Jeffrey Ovall and Gabriel Pinochet-Soto
(Date: January 9, 2025)
1. Introduction
This paper concerns the efficient approximation of several eigenpairs of a class of selfadjoint elliptic operators via an adaptive finite element scheme. We consider eigenvalue problems of the form
(1) $\mathcal{L}\psi=\lambda\psi\mbox{ in }\Omega\;,\;\psi=0\mbox{ on }\partial\Omega\quad,\mbox{ where }\mathcal{L}w=-\nabla\cdot(A\nabla w)+Vw~{},$
and $\Omega\subset\mathbb{R}^{2}$ is a bounded domain. The potential $V$ and matrix $A$ are assumed to be piecewise smooth on some partition of the domain, $\overline{\Omega}=\bigcup_{k=1}^{m}\overline{\Omega_{k}}$ and $\Omega_{i}\cap\Omega_{j}=\emptyset$ when $i\neq j$. We further assume that $V\geq 0$ and that $A$ is uniformly positive definite a.e. on $\Omega$. Since we will discretize the problem using standard finite elements, it is convenient to assume that $\Omega$ and the subdomains $\Omega_{k}$ are polygons, and that all triangulations of $\Omega$ respect the partition (i.e. each triangle is contained in precisely one of the $\Omega_{k}$). The approach we will propose extends naturally to $\mathbb{R}^{3}$, but our numerical experiments will be in 2D, so we focus on that case in the discussion.
Adaptive finite element schemes for eigenvalue computations have a long history, with contributions typically mirroring (and often following soon after) their counterparts for source problems. In the case of computing clusters of eigenvalues, where repeated eigenvalues (or merely tightly clustered eigenvalues) may be present, together with bases for their corresponding invariant subspaces is more a challenging task, and contributions in this direction is relatively recent in comparison with those that concern a few simple eigenvalues near the bottom of the spectrum. We do not attempt a comprehensive survey of the extensive literature on the topic, and instead provide some references that more specifically concern the approximation of several, possibly clustered, eigenvalues, together with a basis of the corresponding invariant subspace [23, 5, 14, 7, 8, 16, 20, 28]. A common theme of nearly all of these contributions it that, whatever strategy one ultimately chooses for adaptively enriching the finite element space used to approximate the eigenpairs, the resulting finite element space should be well-suited to approximate generic functions in the entire invariant subspace. An exception to this pattern is the work [20], in which multiple meshes/FEM spaces are employed.
The generic template for adaptive finite element procedures, for source problems or eigenvalue problems, is to repeat the the following steps until some convergence criterion is met:
SOLVE $\longrightarrow$ ESTIMATE $\longrightarrow$ MARK $\longrightarrow$ REFINE
Global error estimates and local error indicators are obtained in the ESTIMATE step from the approximation(s) obtained in the SOLVE step. In the MARK step, some elements are flagged for refinement based on the local indicators. These elements are refined, either by subdividing the element ($h$ refinement) or increasing the polynomial degree in that element ($p$ refinement). Some other elements, typically near the marked elements, may also need to be refined to maintain some type of conformity. In the case of eigenvalue problems in which multiple eigenpairs are sought, elements are generally marked for refinement based on local error indicators that consider all of the current computed eigenpairs—different approaches primarily differ in how this information is used for marking. A distinguishing feature of the present contribution is that we propose a marking and refinement strategy that is based solely on the approximate solution of a single source problem, namely
(2) $\mathcal{L}u=1\mbox{ in }\Omega\;,\;u=0\mbox{ on }\partial\Omega~{}.$
We will refer to $u$ as the landscape function for the operator and domain, which is what this function is typically called in the context of eigenvector localization (cf. [13, 4, 2, 3]).
Remark 1.1 (Some caveats).
It is possible to construct examples, particularly in 1D, for which the landscape function $u$ is a polynomial (of low degree) in some subdomain $R$, but for which the eigenvectors $\psi_{j}$ of interest are not polynomial on $R$. For such an example we might rightly expect refinement driven by approximations of $u$ to cease refining within $R$ at some stage, and thereby fail to provide convergent eigenvector approximations. Such situations are highly unlikely in practice, however, and we will provide both theory substantial empirical evidence that “landscape refinement” is a very effective approach for the efficient approximation of (large) clusters of eigenvalues and eigenvectors. In Example 4.5, we illustrate a different situation in which a few eigenvectors within a large cluster are converging more slowly under landscape refinement than the rest, and we indicate how such eigenvectors can be detected, extracted, and dealt with efficiently.
The rest of the paper is structured as follows. In Section 2, we provide theoretical results that link the landscape function $u$ to eigenpairs $(\lambda_{j},\psi_{j})$ in a way that makes the efficient approximation of such eigenvectors on finite element spaces based on meshes generated by landscape refinement more reasonable than it may appear at first glance. Preliminary experiments in this section add to the evidence that landscape refinement might actually be a viable alternative to refinement based on the (approximate) eigenpairs of interest. Corollary 2.2, which concerns the pointwise control of properly scaled eigenvectors by $u$ is, by now, a well-known result (cf. [13, 3, 4, 11]). Our more general result, Theorem 2.1, does not appear to be broadly known—we have not found it in the literature. Apart from the initial proof-of-concept computations in Section 2, our numerical experiments are done using the SIPG variant of $hp$ DG finite element discretizations. These are briefly described in Section 3, together with the error estimators and local error indicators that we use in our experiments. Section 4 contains the bulk of the experiments, in which we illustrate the effectiveness of our approach on a variety of problems making relevant comparisons with more traditional approaches. Details about the algorithms used for these comparisons are given in Section 5, and further enhancements of our proposed algorithm are considered in Section 6. A few concluding remarks are given in Section 7
2. Basic Properties of the Landscape Function
The assumptions on $A$ and $V$ imply that we have the classical maximum principle for the operator $\mathcal{L}$. A straight-forward application of the Maximum Principle yields the following refinement of the standard stability estimate, which involves the landscape function in a very natural way. We provide a brief proof for completeness.
Theorem 2.1 (Pointwise Stability Estimate).
If $v\in C(\overline{\Omega})$ and $\mathcal{L}v\in L^{\infty}(\Omega)$, then
(3) $|v(x)|\leq\|v\|_{L^{\infty}(\partial\Omega)}+\|\mathcal{L}v\|_{L^{\infty}(\Omega)}\,u(x)~{},$
for each $x\in\overline{\Omega}$.
Proof.
Let $w_{\pm}=\pm v-\|\mathcal{L}v\|_{L^{\infty}(\Omega)}\,u(x)$. By construction, we have $\mathcal{L}w_{\pm}\leq 0$ on $\Omega$, and it follows by the Maximum Principle that
$w_{\pm}(x)\leq\|w_{\pm}\|_{L^{\infty}(\partial\Omega)}=\|v\|_{L^{\infty}(\partial\Omega)}~{}.$
Rearranging terms yields (3). ∎
The standard stability result replaces $u(x)$ by a constant $C$ in (3). The inequality (3) is sharp, as can be seen by choosing $v=u$.
An immediate consequence of this stability result for the eigenvalue problem (1) is given below.
Corollary 2.2 (Pointwise Control of Eigenvectors).
Let $(\lambda,\psi)$ be an eigenpair of $\mathcal{L}$. It holds that
(4) $\frac{|\psi(x)|}{\lambda\|\psi\|_{L^{\infty}(\Omega)}}\leq u(x)\mbox{ for all }x\in\overline{\Omega}~{}.$
This result first appeared in [13] (with a different proof), and demonstrates an important connection between the landscape function and eigenvectors lower in the spectrum: where the landscape function is relatively small, such eigenvectors are obliged to be relatively small; and where the landscape function is relatively large, such eigenvectors are permitted to be relatively large. We illustrate Corollary 2.2 in Example 2.3. The landscape function, through this result and many others, provides a surprisingly rich understanding of the phenomenon of eigenvector localization, at least lower in the spectrum.
Example 2.3 (1D Example).
In order to illustrate Corollary 2.2, we consider the 1D problem in which $\Omega=(0,1)$ and $\mathcal{L}=-\frac{d}{dx^{2}}+V$. Here, the potential $V$ is piecewise constant on a uniform partition of $\Omega$ into 32 subintervals of equal length, as pictured in Figure 1. It varies between $5274.4361$ and $98928.04$. The corresponding landscape function is also given in this figure, and its local extrema are highlighted. The landscape function in this case has 13 local maxima. The largest of these is roughly $1.3501\times 10^{-4}$, and it occurs near $x=0.42334$; the smallest is roughly $1.5101\times 10^{-5}$, and it occurs near $x=0.20361$. The landscape theory from [13, 4, 2, 3] predicts that there will be a “ground state” eigenvector associated with each of the peaks of $u$, and that each such ground state is largely concentrated (localized) near the corresponding local maximizer. Such a “ground state” has its peak near the corresponding peak of $u$, and decays rapidly away from this peak, generally in such a way that you cannot see any sign changes—we note that there is only one true ground state, associated with the smallest eigenvalue, and that is the only eigenvector that is purely of one sign on $\Omega$. These assertions are illustrated in Figure 2. In this figure, we show scaled versions of the eigenvectors $\psi_{j}$, for $1\leq j\leq 9$ and $47\leq j\leq 49$. Among these, six correspond to what one might naturally consider a ground state, namely those for $j=1,2,3,6,9$ and $j=48$. The eigenvector $\psi_{48}$ is associated with the smallest of the peaks of $u$, and is the last of the ground states. The eigenvectors $\psi_{4}$ and $\psi_{7}$ might also be considered ground states, despite the fact that we can clearly observe a sign change, because their peaks are very closely related to two peaks of $u$ that are not otherwise “accounted for”, in the sense that, if we did not count them as ground states, we would not have 13 ground states to match with the 13 local maxima of $u$. The five remaining ground states, which are not pictured in Figure 2, are $\{\psi_{10},\psi_{13},\psi_{16},\psi_{19},\psi_{27}\}$.
Figure 1. The potential $V$ and associated landscape function $u$ for Example 2.3. The local maxima of $u$ are highlighted.
Figure 2. Several scaled eigenvectors (black) early in the spectrum, together with $u$ and $-u$ (grey), which act as an envelope of the eigenvectors. The eigenvectors are scaled as indicated in (4). Eigenvectors 1-9 and 47-49 for Example 2.3.
We have seen in this example how the landscape function provides remarkably tight control of the behaviour of ground state eigenvectors. We have also seen that, although (4) holds for any eigenpair of $\mathcal{L}$, the bound is much more meaningful lower in the spectrum, for eigenvectors that are not highly oscillatory. Instead of considering how the landscape function governs the behaviour of eigenvectors low in the spectrum, we might consider how such eigenvectors dictate the shape of the landscape function. This perspective is captured in the following result, whose proof is omitted.
Proposition 2.4 (Fourier Expansion of the Landscape Function).
Let $(\lambda_{n},\psi_{n})$, $n\in\mathbb{N}$, be eigenpairs of $\mathcal{L}$ such that $\{\psi_{n}:\,n\in\mathbb{N}\}$ is an orthonormal Hilbert basis (a Fourier basis) of $L^{2}(\Omega)$. The landscape function has an expansion
(5) $u=\sum_{n\in\mathbb{N}}c_{n}\,\psi_{n}\quad,\quad c_{n}=\left(\int_{\Omega}\psi_{n}\,dx\right)/\lambda_{n}~{}.$
Roughly speaking, we expect the dominant features of $u$ to look like a finite linear combination of ground states lower in the spectrum, because it is these vectors that should have the largest coefficients, $c_{n}$, in a relative sense. This assertion is illustrated in Figure 3.
Figure 3. The landscape function $u$ (grey) for Example 2.3 and its partial Fourier expansion (black) using only the 13 ground states. At right, the partial Fourier expansion of $u$ using the first $50$ eigenvectors.
The discussion above was intended to establish some connections between the eigenvectors of $\mathcal{L}$ and the landscape function $u$, as a first indication that adaptive refinement that is aimed at approximating the landscape function might, in fact, yield finite element spaces in which the eigenvectors are also well approximated. We now provide a first illustration of this approach and its performance on a model problem for which there is no eigenvector localization, but for which adaptive refinement is advisable. In Section 4, we provide strong supporting evidence for the “landscape refinement” approach on variety of problems.
Example 2.5 (L-Shape Domain, First Look).
We here consider the Laplacian $\mathcal{L}=-\Delta$ on the domain $\Omega=(-1,1)\times(-1,1)\setminus[0,1)\times[0,1)$. Most eigenvalues/vectors in this case are not known explicitly, though some highly accurate approximations of a few eigenvalues are given in the literature (cf. [31]). It is known is that some eigenvectors have an $r^{2/3}$-singularity near the origin (cf. [21, 22]). This is true of the ground state eigenvector $\psi_{1}$, for instance. So adaptive refinement is natural in this setting.
We approximate the first five eigenpairs of $\mathcal{L}$, all of which are simple, via an $h$-adaptive conforming finite element method on triangular meshes, with fixed local polynomial degree $p=2$. Without elaborating on the details of local error indicators and marking and refinement strategies (they are standard), we provide the final meshes generated by two different refinement strategies in Figure 4. The first is driven by the computed eigenvectors, and the second is driven by the computed landscape function. Both start from the same quasi-uniform coarse mesh. We use (CR) and (LR) respectively to denote “cluster refinement” and “landscape refinement”.
The reference values
$\lambda_{1}=9.639723844\;,\;\lambda_{2}=15.19725193\;,\;\lambda_{3}=19.73920880$
$\lambda_{4}=29.52148111\;,\;\lambda_{5}=31.91263596~{},$
which were computed by the method of particular solutions [6], and are correct in all digits shown, were used to obtain the errors in our eigenvalue approximations. We note that $\lambda_{3}=2\pi^{2}$ and $\psi_{3}=\sin(\pi x)\sin(\pi y)$ (up to scaling). The two final meshes are qualitatively very different, but the eigenvalue approximation errors are comparable; see Figure 2.5. In both cases, the errors $e_{j}=|\hat{\lambda}_{j}-\lambda_{j}|$ in the approximations $\hat{\lambda}_{j}$ exhibit the optimal convergence rates, namely $\mathcal{O}(\mathrm{DOF}^{-2})$. The eigenvalue errors on the finest meshes are
DOF $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
CR 118240 1.38e-07 1.86e-07 3.64e-07 1.04e-06 1.37e-06
LR 119586 1.81e-07 4.19e-07 1.32e-06 6.11e-06 4.32e-06
Although CR wins in each head-to-head comparison of error, the errors are comparable and decrease at the correct rate in both cases, and LR is computationally cheaper, so LR might be considered a viable alternative. In later experiments, in which larger collection of eigenvalues are desired, we will see that LR becomes an even more attractive alternative.
Figure 4. Example 2.5. Top row: two meshes having roughly the same number of degrees of freedom, obtained by CR (left) and LR for the first five eigenvalues. Bottom row: Error plots for CR (left) and LR.
We next consider a cluster of eigenpairs a bit higher in the spectrum. The interval $[50\pi^{2}-10,50^{2}\pi+10]$ contains seven eigenvalues of $\mathcal{L}$; four are simple, but $\lambda=50\pi^{2}$ has a three-dimensional eigenspace
$S=\mathrm{span}\{\sin(7\pi x)\sin(\pi y)\,,\,\sin(5\pi x)\sin(5\pi y)\,,\,\sin(\pi x)\sin(7\pi y)\}~{}.$
Reference values for these eigenvalues, obtained via the method of particular solutions as before, are
$\lambda_{101}$ $=485.71752463708\;,\;\lambda_{102}=490.15998172598\;,\;\lambda_{103,104,105}=493.48022005447$
$\lambda_{106}$ $=499.24106145290\;,\;\lambda_{107}=502.30119419396$
In this case, we consider three different refinement strategies: landscape refinement (LR), eigenvector refinement (CR), and a mixed strategy (MR) that begins with LR and then shifts to CR when the dimension of the finite element space reaches some threshold. The plots in Figure 5 contain the computed eigenvalues in the interval $[50\pi^{2}-10,50^{2}\pi+10]$ for both the LR and CR refinement strategies, starting from the same coarse mesh as before. Early in the adaptive cycle, neither approach has produced a finite element space that is sufficient to resolve the invariant subspace of interest—for the first several refinement cycles, neither approach even computes the correct the number of eigenvalues in the interval interest! In the case of CR, this means that refinement during this phase is being at least partially driven by approximations of that are, at best, related to eigenvectors of $\mathcal{L}$ for eigenvalues outside of $[50\pi^{2}-10,50^{2}\pi+10]$. If CR is generating useful meshes during this phase, it essentially doing so “on accident”. In fact, CR does seem to be generating useful meshes for approximation the invariant subspace of interest, and it does so a bit earlier than LR refinement, based on when we see a discernable pattern of convergence toward the correct eigenvalues; for CR this happens somewhere around 3000 or 4000 DOF, and for LR it is nearer to 8000 DOF.
Figure 5. Example 2.5. Computed approximations of the (seven) eigenvalues in $[50\pi^{2}-10,50\pi^{2}+10]$ over a sequence of adaptively refined meshes using CR (left) and LR. Note that $\lambda_{103}=\lambda_{104}=\lambda_{105}=50\pi^{2}$.
Table 1 below shows the error, $e_{j}=|\lambda_{j}-\hat{\lambda}_{j}|$, in the seven computed eigenvalues $\hat{\lambda}_{j}$ for the first mesh exceeding 100,000 DOF for each of the three refinement strategies. For MR, we show the results of three variants: MR1 switches from LR to CR as soon as DOF $>4000$; MR2 switches when DOF $>8000$; MR3 switches when DOF $>10000$. We first note that, although LR yields the largest final errors among the five options all of these approaches exhibit convergence consistent with the optimal rate $\mathcal{O}(\mathrm{DOF}^{-2})$, despite the fact that the meshes used here for LR are precisely those used for the smallest five eigenvalues. We also remark that, although Figure 5 indicates that CR begins to accurately resolve the eigenvalues a bit sooner than LR, the mixed approach MR3 that transitions from LR to CR latest actually performs best.
Table 1. Example 2.5. Errors $e_{j}=|\lambda_{j}-\hat{\lambda}_{j}|$ on the finest mesh for approximating the seven eigenvalues in $[50\pi^{2}-10,50^{2}\pi+10]$ using several adaptive refinement strategies.
DOF $e_{101}$ $e_{102}$ $e_{103}$ $e_{104}$ $e_{105}$ $e_{106}$ $e_{107}$
CR 110123 4.30e-03 4.29e-03 4.02e-03 4.51e-03 5.11e-03 5.27e-03 5.18e-03
LR 106638 2.57e-02 2.40e-02 2.36e-02 2.42e-02 2.70e-02 2.61e-02 2.30e-02
MR1 104095 4.81e-03 4.70e-03 4.37e-03 4.99e-03 5.67e-03 5.76e-03 5.77e-03
MR2 115258 3.81e-03 3.77e-03 3.57e-03 4.03e-03 4.67e-03 4.67e-03 4.76e-03
MR3 104432 1.96e-03 2.07e-03 1.92e-03 2.06e-03 3.12e-03 2.42e-03 2.59e-03
3. Discontinuous Galerkin discretization
We will be using the $hp$-version SIPG finite element method [1, 19, 17] for the discretization of (1) in Section 4, so we describe it in this section, together with error estimators and the local error indicators that will be used to drive our adaptive algorithms. Throughout, we assume that the computational domain $\Omega$ can be partitioned into a shape-regular mesh $\mathcal{T}$, and that the mesh elements are affine quadrilaterals. The diameter of an element $K$ is denoted by $h_{K}$. In order to be able to deal with irregular meshes we need to define the edges of a mesh ${\mathcal{T}}$. We refer to $e$ as an interior mesh edge of ${\mathcal{T}}$ if $e=\partial K\cap\partial K^{\prime}$ for two neighboring elements $K,K^{\prime}\in{\mathcal{T}}$ whose intersection has a positive surface measure. The set of all interior mesh edges is denoted by $\mathcal{E}_{I}(\mathcal{T})$. Analogously, if the intersection $e=\partial K\cap\partial\Omega$ of the boundary of an element $K\in{\mathcal{T}}$, we refer to $e$ as a boundary mesh edge of ${\mathcal{T}}$. The set of all boundary mesh edges of ${\mathcal{T}}$ is denoted by $\mathcal{E}_{\Gamma}(\mathcal{T})$ and we set $\mathcal{E}(\mathcal{T})=\mathcal{E}_{I}(\mathcal{T})\cup\mathcal{E}_{\Gamma}(\mathcal{T})$. The diameter of an edge $e$ is denoted by $h_{e}$. We allow for 1-irregularly refined meshes $\mathcal{T}$, i.e. we allow for at most one hanging node per edge. We also assume that the discontinuities of $A$ and $V$ are aligned with the edges of the elements.
Next, let us define the jumps and averages of piecewise smooth functions across edges of the mesh ${\mathcal{T}}$. To that end, let the interior edge $e\in{\mathcal{E}}_{I}({\mathcal{T}})$ be shared by two neighboring elements $K_{1}$ and $K_{2}$. For a piecewise smooth function $v$, we denote by $v|_{1,e}$ the trace on $e$ taken from inside $K_{1}$, and by $v|_{2,e}$ the one taken from inside $K_{2}$. The jump of $v$ across the edge $e$ is then defined as
$[\![v]\!]=v|_{1,e}\,\,\underline{n}_{K_{1}}+v|_{2,e}\,\,\underline{n}_{K_{2}}.$
Here, $\underline{n}_{K_{1}}$ and $\underline{n}_{K_{2}}$ denote the unit outward normal vectors on the boundary of elements $K_{1}$ and $K_{2}$, respectively. The average across $e$ of $A\nabla v$ is defined as in [12] for the scalar case:
$\{\!\!\{A\nabla v\}\!\!\}=\omega_{2}(A\nabla v)|_{2,e}+\omega_{1}(A\nabla v)|_{1,e},$
where
$\omega_{2}=\frac{\underline{n}_{K_{2}}^{t}A|_{1,e}\underline{n}_{K_{2}}}{\underline{n}_{K_{2}}^{t}A|_{2,e}\underline{n}_{K_{2}}+\underline{n}_{K_{2}}^{t}A|_{1,e}\underline{n}_{K_{2}}}\ ,\quad\omega_{1}=\frac{\underline{n}_{K_{1}}^{t}A|_{2,e}\underline{n}_{K_{1}}}{\underline{n}_{K_{1}}^{t}A|_{2,e}\underline{n}_{K_{1}}+\underline{n}_{K_{1}}^{t}A|_{1,e}\underline{n}_{K_{1}}}\ .$
On a boundary edge $e\in{\mathcal{E}}_{\Gamma}({\mathcal{T}})$, we accordingly set $\{\!\!\{A\nabla v\}\!\!\}=A\nabla v$ and $[\![v]\!]=v\underline{n}$, with $\underline{n}$ denoting the unit outward normal vector on $\Gamma$.
On each element $K$, we construct a polynomial space of order $p$ like:
(6) $\mathcal{Q}_{p}(K)=\{\,v\,:\,K\to\mathbb{R}\,:\,v\circ T_{K}\in\mathcal{Q}_{p}(\widehat{K})\,\},$
with $\mathcal{Q}_{p}(\widehat{K})$ denoting the set of tensor product polynomials on the reference element $\widehat{K}$ of degree less than or equal to $p$ in each coordinate direction on $\widehat{K}$. In the $hp$- setting, elements in the same mesh might be associated to polynomial spaces of different orders. We assume the order of polynomials to be of bounded local variation, that is, there is a constant $\varrho\geq 1$, independent of the mesh $\mathcal{T}$, such that
(7) $\varrho^{-1}\leq p_{K}/p_{K^{\prime}}\leq\varrho$
where $p_{K}$ and $p_{K^{\prime}}$ are the orders of the elements $K$ and $K^{\prime}$, for any pair of neighboring elements $K,K^{\prime}\in\mathcal{T}$. For a mesh edge $F\in\mathcal{E}(\mathcal{T})$, we introduce the edge polynomial degree $p_{e}$ by
(8) $p_{e}$ =$=\begin{cases}\ \max\{p_{K_{1}},p_{K_{2}}\},&\qquad\text{if }e=\partial K_{1}\cap\partial K_{2}\in{\mathcal{E}}_{I}({\mathcal{T}}),\\[2.84544pt] \ p_{K},&\qquad\text{if }e=\partial K\cap\partial\Omega\in\mathcal{E}_{B}(\mathcal{T}).\end{cases}$
We also denote by $h_{e}$ the length of the edge $e$. For a partition $\mathcal{T}$ of $\Omega$, we define the $hp$-version DG finite element space by
(9) $S(\mathcal{T})=\{\,v\in L^{2}(\Omega)\,:\,v|_{K}\in\mathcal{Q}_{p_{K}}(K),\ K\in{\mathcal{T}}\,\}.$
The standard $L^{2}$ norm on the domain $\Omega$ is denoted by $\|\cdot\|_{0,\Omega}$ and when restricted to an element or edge is denoted by $\|\cdot\|_{0,K}$ and $\|\cdot\|_{0,e}$, respectively. We shall also need the following DG norm:
Definition 3.1 (DG norm).
For any $u\in S(\mathcal{T})$
(10) $\|\,{u}\,\|_{S(\mathcal{T})}^{2}=\sum_{K\in\mathcal{T}}\|A\nabla u-Vu\|^{2}_{0,K}+\sum_{e\in\mathcal{E}({\mathcal{T}})}\frac{\gamma cp_{e}^{2}}{h_{e}}\|[\![u]\!]\|^{2}_{0,e},$
where
$c=2\frac{\underline{n}_{K_{1}}^{t}A|_{1,e}A|_{2,e}\underline{n}_{K_{1}}}{\underline{n}_{K_{1}}^{t}A|_{1,e}\underline{n}_{K_{1}}+\underline{n}_{K_{1}}^{t}A|_{2,e}\underline{n}_{K_{1}}},$
with $A|_{1,e}$ and $A|_{2,e}$ the values of $A$ in the two elements $K_{1}$ and $K_{2}$ sharing $e$. If $e$ is on the boundary, $c=\underline{n}^{t}A\underline{n}$. The penalty parameter $\gamma>0$ appears in the SIPG discretization described below.
The SIPG discrete version of the eigenvalue problem (1) is: find $(\lambda_{h},\phi_{h})\in\mathbb{R}\times S(\mathcal{T})$ such that
(11) $a(\phi_{h},v_{h})=\lambda_{h}\ b(\phi_{h},v_{h})\qquad\forall\,v_{h}\in S(\mathcal{T}),$
where $b(\cdot,\cdot)$ is the $L^{2}$ inner product and with $\|\phi_{h}\|_{0,\Omega}=1$. The bilinear form $a(u,v)$ is given by
[…]
where the gradient operator $\nabla$ is defined elementwise and the parameter $\gamma>0$ is the interior penalty parameter. For source problems the SIPG is known to be a stable and consistent method for sufficiently large values of $\gamma$ [1, 29].
The SIPG discrete version of the landscape source problem (2) is: find $u_{h}\in\mathbb{R}\times S(\mathcal{T})$ such that
(13) $a(u_{h},v_{h})=b(1,v_{h})\qquad\forall\,v_{h}\in S(\mathcal{T}).$
Two residual-based error estimators are used in the simulations, one for the landscape solution $u$ (2) which is based on the error estimator for linear problems, and one for the eigenpairs of the eigenvalue problem (1). An a posteriori error estimator is a necessary ingredient for adapting the mesh and the finite element space because it is capable of estimating the error in each element of the computed solution using only the computed solution and given data for the problem.
The error estimator for the landscape function $\eta_{\mathrm{land}}$ is based on [24, 15] and is computed using an approximation $u_{h}$ of $u$ computed on $S(\mathcal{T})$:
(14) $\eta_{\mathrm{land}}^{2}:=\sum_{K\in\mathcal{T}}\eta_{\mathrm{land},K}^{2}\;,\;\eta_{\mathrm{land},K}^{2}:=\eta_{R_{\mathrm{land},K}}^{2}(u_{h})+\eta_{F_{K}}^{2}(u_{h})+\eta_{J_{K}}^{2}(u_{h})$
where
$\eta_{R_{\mathrm{land},K}}^{2}(u_{h})\ :=$ $\ A_{\mathrm{min},K}^{-1}p_{K}^{-2}h_{K}^{2}\|1+\nabla\cdot(A\nabla u_{h})-Vu_{h}\|^{2}_{0,K}\ ,$
$\eta_{F_{K}}^{2}(u_{h}):=$ $\frac{1}{2}\sum_{e\in\mathcal{E}_{I}(\mathcal{T})}A_{\mathrm{min},e}^{-1}\,p_{e}^{-1}h_{e}\|[\![A\nabla u_{h}]\!]\|^{2}_{0,e}\ ,$
$\eta_{J_{K}}^{2}(u_{h})\ :=$ $\frac{1}{2}\sum_{e\in\mathcal{E}_{I}(\mathcal{T})}\Bigg{(}\frac{h_{e}}{A_{\mathrm{min},e}\,p_{e}}+\frac{\gamma^{2}A_{\mathrm{max},e}\,p_{e}^{3}}{h_{e}}\Bigg{)}\|[\![u_{h}]\!]\|^{2}_{0,e}$
$+\sum_{e\in\mathcal{E}_{\Gamma}(\mathcal{T})}\Bigg{(}\frac{h_{e}}{A_{\mathrm{min},e}\,p_{e}}+\frac{\gamma^{2}A_{\mathrm{max},e}\,p_{e}^{3}}{h_{e}}\Bigg{)}\|u_{h}\|^{2}_{0,e}\ ,$
where $A_{\mathrm{min},e}$ and $A_{\mathrm{max},e}$ are respectively the minimum and the maximum eigenvalue of $A$ among the elements sharing the edge $e$ and where $A_{\mathrm{min},K}$ is the minimum eigenvalue of $A$ in $K$. When $A\equiv 1$ everywhere, the term $\frac{h_{e}}{p_{e}}\lesssim\frac{p_{e}^{3}}{h_{e}}$ in $\eta_{J_{K}}^{2}(u_{h})$. Therefore, the term $\eta_{J_{K}}^{2}(u_{h})$ reduces to:
$\eta_{J_{K}}^{2}(u_{h})\ :=$ $\frac{1}{2}\sum_{e\in\mathcal{E}_{I}(\mathcal{T})}\frac{\gamma^{2}p_{e}^{3}}{h_{e}}\|[\![u_{h}]\!]\|^{2}_{0,e}+\sum_{e\in\mathcal{E}_{\Gamma}(\mathcal{T})}\frac{\gamma^{2}p_{e}^{3}}{h_{e}}\|u_{h}\|^{2}_{0,e}\ .$
From [24] we have that $\eta_{\mathrm{land}}$ is reliable:
(15) $\|u-u_{h}\|_{S(\mathcal{T})}\lesssim\eta_{\mathrm{land}}\ .$
The error estimator $\eta_{\mathrm{eig},j}$ for the eigenpairs of the eigenvalue problem (1) is based on [18, 19] and computed for each eigenpair $(\lambda_{j},\phi_{j})$ using an approximation $(\lambda_{j,h},\phi_{j,h})$ of the eigenpair computed on $S(\mathcal{T})$:
(16) $\eta_{\mathrm{eig},j}^{2}:=\sum_{K\in\mathcal{T}}\eta_{\mathrm{eig},K,j}^{2}\;,\;\eta_{\mathrm{eig},K,j}^{2}=\eta_{R_{\mathrm{eig},K}}^{2}(\lambda_{j,h},\phi_{j,h})+\eta_{F_{K}}^{2}(\phi_{j,h})+\eta_{J_{K}}^{2}(\phi_{j,h})$
where
$\eta_{R_{\mathrm{eig},K}}^{2}(\lambda_{j,h},\phi_{j,h})\ :=\ A_{\mathrm{min},K}^{-1}p_{K}^{-2}h_{K}^{2}\|\lambda_{j,h}\phi_{j,h}+\nabla\cdot(A\nabla\phi_{j,h})-V\phi_{j,h}\|^{2}_{0,K}\ .$
The “jump terms” $\eta_{F_{K}}$ and $\eta_{J_{K}}$ here are defined in the same way they were for the landscape function. From [19] we have that $\eta_{\mathrm{eig},j}$ is reliable for both eigenvalues and eigenvectors:
(17) $|\lambda_{j}-\lambda_{j,h}|\lesssim\eta_{\mathrm{eig},j}^{2}+\mathrm{h.o.t.}\ ,$
(18) $\mathrm{dist}(\phi_{j,h},E_{1}(\lambda_{j}))\lesssim\eta_{\mathrm{eig},j}+\mathrm{h.o.t.}\ ,$
where $\mathrm{h.o.t.}$ are asymptotically higher-order terms that are not numerically computed or approximated and where $E_{1}(\lambda_{j})$ is the span of all eigenfunctions of the eigenvalue $\lambda_{j}$ normalized in the $L^{2}$ norm and where $\mathrm{dist}(\phi_{j,h},E_{1}(\lambda_{j}))$ is the distance in the DG norm of the computed eigenvector $\phi_{j,h}$ from the eigenspace $E_{1}(\lambda_{j})$.
4. Numerical Results
We will compare landscape refinement with versions of eigenvector-based refinement on several examples having different features. We first define of all the quantities used in presenting the results that are collected, starting from the quantities related to the eigenvalue problem:
  • Error bound for single eigenpair: for the computed eigenpair $(\lambda_{j,h},\phi_{j,h})$ of index $j$, the error bound is $\eta_{\mathrm{eig},j}^{2}$.
  • Relative error bound for single eigenpair: for the computed eigenpair $(\lambda_{j,h},\phi_{j,h})$ of index $j$, the error bound is $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$.
  • Envelope for the true relative error for eigenvalues: This quantity is only computable if the true values of the eigenvalues are available. For an approximation of the lowest $M$ eigenpairs in the spectrum, it is defined as $e_{\mathrm{max}}^{\mathrm{rel}}=\max_{j\leq M}(|\lambda_{j}-\lambda_{j,h}|/\lambda_{j})$.
  • Envelope for the relative error bound for eigenvalues: for an approximation of the lowest $M$ eigenpairs in the spectrum, it is defined as $\eta_{\mathrm{max},\mathrm{rel}}^{2}=\max_{j\leq M}(\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h})$.
The only quantity for the landscape problem is the error bound for the landscape solution defined as $\eta_{\mathrm{land}}$.
The last quantity to introduce is CPU time which is the sum of the time spent by all processing units on a task. For example, if two cores of a modern CPU are working at the same time on a task for 1 minute then the CPU time for the task is 2 minutes. CPU time is a better measure of the computational complexity of a task than real time because it is independent of the number of cores used. Using CPU time different codes with different levels of parallelization can be compared in a meaningful way.
Although we will primarily compare landscape refinement with cluster refinement, there will be a few cases in which we also consider refinement driven by a single eigenpair. We will refer to this as eigenpair refinement. In order to present the experiments sooner, details of the particular algorithms will be provided in Section 5. Here, we merely identify the highest-level algorithms with the descriptors above:
  1. (1)
    Eigenpair refinement (ER) for a single eigenpair, Algorithm 1
  2. (2)
    Cluster refinement (CR) based on the first $M$ eigenpairs, Algorithm 2
  3. (3)
    Landscape refinement (LR), Algorithm 3
Algorithms 2 and 3 use the largest of the eigenvalue error estimates for the cluster to determine the stopping criterion. The key difference between Algorithms 2 and 3 is the strategy for marking elements and deciding how to refine them, with Algorithm 3 using the landscape function for these decisions and Algorithm 2 using some collective information from the cluster.
Example 4.1 (Laplace operator on the unit square).
We consider the operator $\mathcal{L}w=-\Delta w$ on $\Omega=(0,1)^{2}$, for which the pairs are well-known, $\lambda_{mn}=(m\pi)^{2}+(n\pi)^{2}$ and $\phi_{mn}=2\sin(m\pi x)\sin(n\pi y)$ for $m,n\in\mathbb{N}$. The eigenvectors are analytic and do not localize anywhere, but we will see that landscape refinement is effective even in such simple situations.
In Figure 6, we show the collective eigenvalue error and error estimate, $e_{\mathrm{max}}^{\mathrm{rel}}$ and $\eta_{\mathrm{max},\mathrm{rel}}^{2}$, for the first $M=100$ eigenpairs, under landscape refinement starting from an initial mesh of $8\times 8$ elements and initial polynomial order $p=2$ in each element. This figure also includes relative error bounds for some individual eigenvalues in the cluster. For comparison, the error bound for the landscape solution $\eta_{\mathrm{land}}^{2}$ is also shown. The reliability of the underlying eigenvector error estimator (17) is reflected in the very similar convergence patters of $e_{\mathrm{max}}^{\mathrm{rel}}$ and $\eta_{\mathrm{max},\mathrm{rel}}^{2}$. The square of the error bound for the landscape solution $\eta_{\mathrm{land}}^{2}$ converges faster and this is because the curves for envelopes are dominated by the error from the eigenpairs in the higher part of the computed spectrum. Such eigenpairs have high frequency and therefore they are harder to approximate in FEMs. On the other hand, the landscape solution has a very low frequency so it converges faster. It is remarkable that adapting for the landscape solution, which has low frequency, causes high frequency eigenpairs to converge.
Figure 6. Example 4.1 Convergence plots of collective and individual eigenvalue errors under landscape refinement. Left: The square of the error bound for the landscape solution $\eta_{\mathrm{land}}^{2}$ and the envelopes for the relative error bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ and the true error for the computed spectrum $e_{\mathrm{max}}^{\mathrm{rel}}$. Right: Relative errors for some individual eigenvalues.
Figure 7 contains convergence comparisons between landscape refinement, eigenvector refinement based on single eigenpairs, and two variants of cluster refinement for the first $M=100$ eigenpairs. The two variants involve how local error and regularity indicators for each eigenpair in the cluster are used for purposes of adaptivity. The two approaches, AdaptSpectrumEigSum and AdaptSpectrumEigMax, are given in detail in Section 5 as Algorithm 7 and Algorithm 8, respectively, but the names themselves are suggestive of how they operate.
The first picture in Figure 7 demonstrates that landscape refinement can beat, or at least be competitive with, eigenvector refinement for driving down the error in a single eigenvalue. We run eigenvector refinement independently three times, targeting the eigenpairs of indices 1, 48 and 90. These convergence histories are plotted together with the convergence for these same eigenvalues computed on meshes determined by landscape refinement. For the first eigenvalue, eigenpair refinement leads to faster convergence than landscape refinement, but both are extremely fast. What is more interesting is that landscape refinement outperforms eigenpair refinement for the other eigenvalues. More specifically, after an initial phase in which both approaches yield similar convergence behavior, the landscape approach yields two significant drops in estimated eigenvalue error between $40^{2}$ and $60^{2}$ DOFs that mimic similar drops in the estimated (square) landscape error. It may seem counterintuitive that landscape refinement can beat eigenpair refinement for single eigenpairs, but perhaps it is not so surprising in this case because, on coarse meshes, the approximate eigenvectors having index 48 and 90 are not approximating the actual 48th and 90th eigenvectors well (if at all), so the refinement dictated at those levels is based on poor information. In contrast, the landscape function is well-resolved even on coarse meshes, and provides accurate information early on concerning where and how to refine the mesh in a way that is useful even for the higher-frequency eigenvectors (though it is a low-frequency function). The second picture in Figure 7 shows that landscape refinement convincingly beats both variants of cluster refinement. Based on what we observed in the single eigenvector refinement cases, this is not a surprise, though it merits reporting.
Figure 7. Example 4.1. Left: Convergence plots of the relative eigenvalue errors $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ under landscape refinement and single eigenpair refinement for $j=1,48,90$. Right: Convergence plots of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ under landscape refinement two variants of cluster refinement. Also included is the error bound for the landscape function $\eta_{\mathrm{land}}^{2}$ under landscape refinement.
Example 4.2 (Laplace operator on the L-shape domain, revisited).
We briefly revisit Example 2.5, but with $hp$ adaptivity using landscape refinement and cluster refinement for the first $M=100$ eigenpairs. As before, we start with initial mesh of $8\times 8$ elements of order $p=2$. The convergence results for landscape refinement and cluster refinement using AdaptSpectrumEigSum are shown in Figure 8, for the envelope for the relative error bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$. The final meshes in each case are given in Figure 9. We see that landscape refinement, though initially providing slightly slower convergence, overtakes cluster refinement at around $140^{2}$ DOFs, and reaches an error tolerance of $10^{-4}$ for the cluster much(!) sooner. The final mesh in the case of landscape refinement is also more appealing, with $h$-refinement targeted near the corners of the domain, whereas the final mesh for cluster refinement is far less coherent.
Figure 8. Example 4.2. Convergence plots of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ under landscape refinement two variants of cluster refinement. Also included is the error bound for the landscape function $\eta_{\mathrm{land}}^{2}$ under landscape refinement.
Figure 9. Example 4.2. Final mesh computed using landscape refinement (top) and cluster refinement (bottom). Colors represent local polynomial degree.
We briefly comment on the high polynomial degrees seen in the corners of the landscape-adapted mesh in Figure 9, which may seem counter-intuitive as the elements are small there. A wealth of empirical evidence indicates that such seeing high polynomial degrees in regions with small elements is not at all unusual for many $hp$-adaptive schemes, and does not negatively impact the exponential convergence rate. We note the works [10, 9] for their use of $h$ and/or $p$ coarsening during the adaptive process. Under such a scheme, we might not have obtained high polynomial degrees near the corners, but the exponential convergence would not have been noticeably improved.
Example 4.3 (Simple Schrödinger operator).
Here we consider the Schrödinger operator $\mathcal{L}w=-\Delta w+Vw$ on the domain $\Omega=(0,1)^{2}$, where $V$ is the piecewise constant potential defined in Figure 10, whose values range between 0 and 6400. Again, we start with an initial mesh of $8\times 8$ elements and with initial order $p=2$, and consider landscape refinement, single eigenpair refinement, and cluster refinement for the first $M=100$ eigenpairs. In the case of cluster refinement, we use AdaptSpectrumEigSum for marking elements and assessing local regularity.
In Figure 10, we also see the convergence of the relative error bound $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ for the $j$th (single) eigenpair under each of the three refinement schemes. Concerning the comparison between landscape refinement and eigenpair refinement, we see similar behavior as we did in Figure 6 for the Laplacian on this domain. The convergence for $j=1$ was slightly better under eigenpair refinement than for landscape refinement, reaching $10^{-4}$ first, and cluster refinement was the worst. Higher in the spectrum, $j=57$ and $j=76$, eigenpair refinement and cluster refinement performed similarly, beating landscape refinement until around $200^{2}$ DOF, when landscape refinement began to significantly overtake them.
Figure 10. Example 4.3. Top left: The potential $V$. Other plots: Convergence of the relative error bound $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ for the $j$th eigenpair under landscape refinement, cluster refinement for the first $M=100$ eigenpairs, and single eigenvector refinement for the targeted eigenpair, $j=1,57,76$.
Example 4.4 (Schrödinger Operator with Rough Potential).
We again take $\Omega=(0,1)^{2}$ and the Schrödinger operator $\mathcal{L}w=-\Delta w+Vw$, where $V$ is the piecewise constant potential defined in Figure 11, with values ranging between about 19.23 and 7942 on a $20\times 20$ square partition of the domain, similar to the examples presented in [13]. These values were independently drawn from a uniform distribution on $[0,8000]$. Also shown in this figure is the landscape function computed in the finest space generated by landscape refinement. The initial mesh used order $p=2$ elements on the $20\times 20$ mesh used to define $V$. This was the type of problem that originally motivated our consideration of landscape refinement, though we have already seen that landscape refinement performs quite well on other problems for which the landscape function does not exhibit such strong localization behavior.
Also in Figure 11 are convergence plots for the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ for the first $M=100$ eigenpairs, and several individual relative eigenvalue errors $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$, under landscape refinement. For comparison, we also include the square of the error bound for the landscape solution $\eta_{\mathrm{land}}^{2}$. We note that, although the individual eigenvectors have very different behaviors, landscape refinement yields very similar convergence for each of them. Apart from scaling, this convergence tracks that of $\eta_{\mathrm{land}}^{2}$ quite well, i.e. all eigenpairs converge at about the same speed as the landscape solution.
Figure 11. Example 4.4. Top row: The potential $V$ and the associated landscape function. Bottom row: The relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ (left) and several individual relative eigenvalue error bounds $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ (right) under landscape refinement, plotted together with $\eta_{\mathrm{land}}^{2}$.
Example 4.5 (Discontinuous Diffusion).
Here, we consider the operator $\mathcal{L}w=-\nabla\cdot(A\nabla w)$, where $A$ is piecewise constant with value 1 in $(0,0.5]^{2}$ and $[0.5,1)^{2}$ and $\beta^{2}=10$ elsewhere. We consider the first $M=50$ eigenpairs. This kind of discontinuous diffusion was considered by Kellogg [26] for the source problem, and the results there suggest that some eigenvectors be singular near $(1/2,1/2)$, with asymptotic behavior like $r^{\alpha}$ for some $\alpha\in(0,1)$, where $r$ is the distance to $(1/2,1/2)$. In this case, $\alpha$ can be as small as $\alpha=(4/\pi)\mathrm{arccot}(\beta)\approx 0.389964$; see [18], where a related problem on the unit disk is worked out analytically.
Convergence of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ and some individual relative error bounds $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ are given for both landscape refinement and cluster refinement, in Figure 12. For cluster refinement, AdaptSpectrumEigSum is used. As before, $\eta_{\mathrm{land}}^{2}$, computed under landscape refinement, is provided for comparison. Again, we begin with order $p=2$ elements on an $8\times 8$ square partition of the domain.
Figure 12. Example 4.5. Top left: Convergence plots of the square of the error bound for the landscape solution $\eta_{\mathrm{land}}^{2}$ and the envelope for the relative error bounds $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ computed using landscape refinement and cluster refinement. Top right: Convergence rates of relative eigenvalue error estimates under landscape refinement for several eigenvalues within the cluster. Bottom left: Convergence rates of eigenvectors 4, 8, 14 and 44 under single eigenpair refinement, superimposed over the top left plot. Bottom right: Range of numerically estimated local regularity measures on the elements of the final mesh for each eigenpair. The landscape function is indicated in red, the slow-converging eigenpairs are in black, and the rest are in blue.
Landscape refinement overtakes cluster refinement near $150^{2}$ DOF, but the convergence curves remain much closer to each other than in previous examples, and the convergence, though still exponential, is slower. We also see in this figure the convergence of several individual relative eigenvalue error estimates within the cluster. The eigenvalues with indices 1, 40 and 50 converge like $\eta_{\mathrm{max},\mathrm{rel}}^{2}$, whereas 4, 8, 14, and 44 converge much more slowly. We highlight, however, that this slower convergence is still exponential.
A likely explanation of this is that the landscape function is more regular (less singular near $(1/2,1/2)$) than these eigenvectors. It is quite possible, due to symmetries in the operator, that the eigenvectors with the strongest singular near the $(1/2,1/2)$ have average value zero, so they do not appear in the Fourier expansion of the landscape function $u$. Based on our computations, this appears to be the case. Figure 12 provides an analysis of the numerically estimated regularity for all eigenpairs in the cluster on the finest mesh. These local regularity measures were computed based on the algorithm described in [25], which uses the decay rates coefficients of local Legendre expansions of the function under consideration to assess regularity. Without getting into detail here, we remark that these local regularity measures are between 00 and $1$, and that larger measures indicate less regularity. Each vertical segment represents the ranges of the regularity measures for the corresponding function. We see that eigenvectors 4, 8, 14, and 44 are among the 11 eigenvectors in this cluster having the lowest regularity (22% of them), and the landscape function $u$ has the highest regularity. Since landscape refinement uses this local regularity indicator to determine how to refine a marked element ($h$ vs. $p$), this makes sense of the slower convergence of the less regular eigenvectors.
Although all eigenvectors are converging exponentially under landscape refinement, one might consider various ways of speeding up the convergence of slower ones. The simplest thing to attempt in this regard would be to decrease the parameter $\mathrm{tol}_{\mathrm{ana}}$ in the landscape refinement algorithm (Algorithm 7) to force $h$-refinement for more of the marked elements. The value of this parameter used in all experiments in this section so far is $\mathrm{tol}_{\mathrm{ana}}=0.25$. We briefly summarize the results of choosing $\mathrm{tol}_{\mathrm{ana}}\in\{0.1,0.001,0.0001\}$: decreasing $\mathrm{tol}_{\mathrm{ana}}$ caused the convergence rates of eigenpairs to get closer to each other, but did not really improve individual or overall convergence rates. A more efficient way to deal with slow converging eigenpairs would be to flag them either by looking at their convergent rates or their local regularity measures, and treating these eigenvectors differently. There are a few natural ways of doing so. Once the slow converging eigenvectors are identified, one might use the worst of the local regularity measures for these eigenvectors and the landscape function to decide how to refine a marked element. Alternatively, one might “extract” them from the cluster and continue their approximation individually on multiple meshes using single eigenpair refinement, or perhaps collectively on a single mesh using cluster refinement for those extracted eigenpairs. We do not develop these variants here, but instead show the effects of single eigenpair refinement on eigenpairs 4, 8, 14 and 44 in Figure 12, overlaid on the convergence plot for the entire cluster under landscape and cluster refinements, to suggest how flagging slow converging eigenpairs under landscape refinement for alternative treatment can be beneficial.
5. The Algorithms
The algorithms are organized as follows:
  1. (1)
    Algorithm 1 (single eigenpair refinement). The marking and refinement strategies are given in Algorithm 4 (AdaptSingleEig).
  2. (2)
    Algorithm 2 (cluster refinement). The marking and refinement strategies are given either by Algorithm 5 (AdaptSpectrumEigSum) or Algorithm 6 (AdaptSpectrumMax).
  3. (3)
    Algorithm 3 (landscape refinement). The marking and refinement strategies are given in Algorithm 7 (AdaptLandscape).
Algorithm 1 is standard, and will not be discussed in any detail here, apart from clarifying that, although it targets the $j$th eigenpair, the first $j$ eigenpairs are computed to ensure that we really have the $j$th in hand to drive the refinement. Both Algorithms 2 and 3 target the first $M$ eigenpairs of the operator, and both use the same criterion for convergence. They differ (significantly) in how they go about marking element for refinement and deciding how to refine them. In the case of Algorithm 3, these decisions are based on the current approximation of the landscape function $u_{h}$, as described in Algorithm 7. For Algorithm 2, all of the computed eigenpairs are used in some way to determine which elements are to be refined, and how, and two natural strategies are given in Algorithms 5 and 6. The two key parameters in Algorithms 4-7 are $r$, which is the percentage of element to be marked for refinement, and $\mathrm{tol}_{\mathrm{ana}}$, which dictates the choice of $h$ versus $p$ refinement for a marked element based on a computed measure of the local regularity of a function in that element. The default settings for these parameters for our experiments, are $r=10$% and $\mathrm{tol}_{\mathrm{ana}}=0.25$. The local regularity measure is computed using the method presented in [25], which is based on the coefficient decay of the local Legendre series expansion. This assessment takes place on line 4 of Algorithm 4, line 6 of Algorithm 5, line 5 of Algorithm 6, and line 4 of Algorithm 7.
Algorithm 1 Adaptive algorithm for a single eigenpair
Input
            Initial FE space $S(\mathcal{T})$.
            Index $j$ of the eigenpair to compute.
            Maximum number of adaptive steps $n_{\mathrm{max}}$.
            Relative tolerance $\mathrm{tol}$.
      Output
            Approximated eigenpair $(\lambda_{h},\phi_{h})$.
$n:=1$
2:while $n<n_{\mathrm{max}}$ do
3:     Compute an approximation of $(\lambda_{j,h},\phi_{j,h})$ computing the first $j$ eigenpairs at the bottom of the spectrum for the problem.
4:     Compute the error estimator $\eta_{\mathrm{eig},j}$.
5:     if $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}<\mathrm{tol}$ then
6:         $(\lambda_{h},\phi_{h}):=(\lambda_{j,h},\phi_{j,h})$ and exit.
7:     else
8:         Adapt $S(\mathcal{T})$ using $\eta_{\mathrm{eig},j}$ and $\phi_{h}$ by calling AdaptSingleEig.
9:     end if
10:     $n:=n+1$
11:end while
Algorithm 2 Adaptive algorithm for a portion of the spectrum
Input
            Initial FE space $S(\mathcal{T})$.
            Index $M$ of the highest eigenpair to compute.
            Maximum number of adaptive steps $n_{\mathrm{max}}$.
            Relative tolerance $\mathrm{tol}$.
      Output
            Approximated eigenpairs $(\lambda_{j,h},\phi_{j,h})$, for all $j=1,\dots,M$.
$n:=1$
2:while $n<n_{\mathrm{max}}$ do
3:     Compute an approximation of $(\lambda_{j,h},\phi_{j,h})$ computing the first $M$ eigenpairs at the bottom of the spectrum for the problem.
4:     Compute the error estimator $\eta_{\mathrm{eig},j}$, for all $j=1,\dots,M$.
5:     if $\max_{j\leq M}(\eta_{\mathrm{eig},j}^{2})<\mathrm{tol}$ then
6:         Return $(\lambda_{j,h},\phi_{j,h})$, for all $j=1,\dots,M$, and and exit.
7:     else
8:         Adapt $S(\mathcal{T})$ using $\eta_{\mathrm{eig},j}$ and $\phi_{j,h}$, for all $j=1,\dots,M$ by calling either AdaptSpectrumEigSum orAdaptSpectrumEigMax.
9:     end if
10:     $n:=n+1$
11:end while
Algorithm 3 Adaptive algorithm using the landscape function
Input
            Initial FE space $S(\mathcal{T})$.
            Index $M$ of the highest eigenpair to compute.
            Maximum number of adaptive steps $n_{\mathrm{max}}$.
            Relative tolerance $\mathrm{tol}$.
      Output
            Approximated eigenpairs $(\lambda_{j,h},\phi_{j,h})$, for all $j=1,\dots,M$.
$n:=1$
2:while $n<n_{\mathrm{max}}$ do
3:     Compute an approximation of $(\lambda_{j,h},\phi_{j,h})$ computing the first $M$ eigenpairs at the bottom of the spectrum for the problem.
4:     Compute an approximation $u_{h}$ of the solution of the landscape function problem.
5:     Compute the error estimator $\eta_{\mathrm{eig},j}$, for all $j=1,\dots,M$.
6:     Compute the error estimator $\eta_{\mathrm{land}}$.
7:     if $\max_{j\leq M}(\eta_{\mathrm{eig},j}^{2})<\mathrm{tol}$ then
8:         Return $(\lambda_{j,h},\phi_{j,h})$, for all $j=1,\dots,M$, and and exit.
9:     else
10:         Adapt $S(\mathcal{T})$ using $\eta_{\mathrm{land}}$ and $u_{h}$ by calling AdaptLandscape.
11:     end if
12:     $n:=n+1$
13:end while
Algorithm 4 AdaptSingleEig: adaptation of the mesh using a single eigenpair
Input
            Current FE space $S(\mathcal{T})$.
            Percentage $r$ of elements to refine.
            Error estimator $\eta_{\mathrm{eig},j}$
            Computed solution $\phi_{h}$.
            Threshold $\mathrm{tol}_{\mathrm{ana}}$ for the smoothness criteria.
      Output
            Adapted FE space $S(\mathcal{T})$.
1:Order the elements of $\mathcal{T}$ in decreasing order with respect to $\eta_{\mathrm{eig},K,j}^{2}$.
2:Mark $r$-percent of the elements from the top of the sorted list. Create a set $\mathcal{M}$ of the marked elements.
3:for all $K\in\mathcal{M}$ do
4:     Estimate the smoothness of $\phi_{h}$ in $K$.
5:     if The smoothness is below $\mathrm{tol}_{\mathrm{ana}}$  then
6:         Mark the element $K$ to be refined in $p$.
7:     else
8:         Mark the element $K$ to be refined in $h$.
9:     end if
10:end for
11:Enforce local properties for marked elements.
12:Adapt $S(\mathcal{T})$.
Algorithm 5 AdaptSpectrumEigSum: adaptation of the mesh using the sum of all computed eigenpairs
Input
            Current FE space $S(\mathcal{T})$.
            Percentage $r$ of elements to refine.
            Error estimators $\eta_{\mathrm{eig},j}$, for all $j=1,\dots,M$
            Computed eigenvectors $\phi_{j,h}$, for all $j=1,\dots,M$.
            Threshold $\mathrm{tol}_{\mathrm{ana}}$ for the smoothness criteria.
      Output
            Adapted FE space $S(\mathcal{T})$.
1:For each element $K$, calculate $\eta_{\mathrm{sum},K}^{2}=\sum_{j=1}^{M}\eta_{\mathrm{eig},K,j}^{2}$.
2:Sum all the computed eigenpairs $\phi_{\mathrm{sum},h}=\sum_{j=1}^{M}\phi_{j,h}$.
3:Order the elements of $\mathcal{T}$ in decreasing order with respect to $\eta_{\mathrm{sum},K}^{2}$.
4:Mark $r$-percent of the elements from the top of the sorted list. Create a set $\mathcal{M}$ of the marked elements.
5:for all $K\in\mathcal{M}$ do
6:     Estimate the smoothness of $\phi_{\mathrm{sum},h}$ in $K$.
7:     if The smoothness is below $\mathrm{tol}_{\mathrm{ana}}$  then
8:         Mark the element $K$ to be refined in $p$.
9:     else
10:         Mark the element $K$ to be refined in $h$.
11:     end if
12:end for
13:Enforce local properties for marked elements.
14:Adapt $S(\mathcal{T})$.
Algorithm 6 AdaptSpectrumEigMax: adaptation of the mesh using the max of all computed eigenpairs
Input
            Current FE space $S(\mathcal{T})$.
            Percentage $r$ of elements to refine.
            Error estimators $\eta_{\mathrm{eig},j}$, for all $j=1,\dots,M$
            Computed eigenvectors $\phi_{j,h}$, for all $j=1,\dots,M$.
            Threshold $\mathrm{tol}_{\mathrm{ana}}$ for the smoothness criteria.
      Output
            Adapted FE space $S(\mathcal{T})$.
1:For each element $K$, calculate $\eta_{\mathrm{max},K}^{2}=\max_{j=1}^{M}\{\eta_{\mathrm{eig},K,j}^{2}\}$.
2:Order the elements of $\mathcal{T}$ in decreasing order with respect to $\eta_{\mathrm{max},K}^{2}$.
3:Mark $r$-percent of the elements from the top of the sorted list. Create a set $\mathcal{M}$ of the marked elements.
4:for all $K\in\mathcal{M}$ do
5:     Estimate the smoothness of $\phi_{i,h}$ in $K$ where $i=\mathrm{argmax}_{j=1}^{M}\{\eta_{\mathrm{eig},K,j}^{2}\}$.
6:     if The smoothness is below $\mathrm{tol}_{\mathrm{ana}}$  then
7:         Mark the element $K$ to be refined in $p$.
8:     else
9:         Mark the element $K$ to be refined in $h$.
10:     end if
11:end for
12:Enforce local properties for marked elements.
13:Adapt $S(\mathcal{T})$.
Algorithm 7 AdaptLandscape: adaptation of the mesh for linear problems
Input
            Current FE space $S(\mathcal{T})$.
            Percentage $r$ of elements to refine.
            Error estimator $\eta_{\mathrm{land}}$
            Computed solution $u_{h}$.
            Threshold $\mathrm{tol}_{\mathrm{ana}}$ for the smoothness criteria.
      Output
            Adapted FE space $S(\mathcal{T})$.
1:Order the elements of $\mathcal{T}$ in decreasing order with respect to $\eta_{\mathrm{land},K}^{2}$.
2:Mark $r$-percent of the elements from the top of the sorted list. Create a set $\mathcal{M}$ of the marked elements.
3:for all $K\in\mathcal{M}$ do
4:     Estimate the smoothness of $u_{h}$ in $K$.
5:     if The smoothness is below $\mathrm{tol}_{\mathrm{ana}}$  then
6:         Mark the element $K$ to be refined in $p$.
7:     else
8:         Mark the element $K$ to be refined in $h$.
9:     end if
10:end for
11:Enforce local properties for marked elements.
12:Adapt $S(\mathcal{T})$.
6. Optimisation of the algorithm
Considering Example 4.3 as a model problem for illustrative purposes, we see in Table 2 that almost 80% of the CPU time is spent on the eigensolver. Moreover, in Figure 13, it is clear that the CPU needed by the eigensolver is growing faster than any other part of the code. In the next two subsections, different ways to optimise and speed up Algorithm 3, the landscape refinement algorithm, are presented.
Algorithm Parts CPU Time (%)
Eigensolver 77.6
Solver 10.7
Error Estimator Eigenproblem 9.2
Error Estimator Landscape 0.2
Other 2.3
Table 2. Cumulative CPU time for different parts of Algorithm 3 on Example 4.1.
Figure 13. CPU time of different parts of the code during the adaptation of the mesh for Algorithm 3.
6.1. Pausing the eigensolver
Since the computed eigenpairs are not used to refine the mesh, but only in the stopping criterion in Algorithm 3, it makes sense to not call the eigensolver in every iteration. Algorithm 8 is a modification of Algorithm 3 where a pause is introduced for the eigensolver. The length of the pause is controlled by $\ell_{\mathrm{pause}}\geq 0$ and measured in the number of iterations. The eigensolver is always called during the first iteration. The algorithm cannot terminate while the eigensolver is paused because the stopping criterion is based on the computed spectrum.
Algorithm 8 Pausing adaptive algorithm using the landscape function
Input
            Initial FE space $S(\mathcal{T})$.
            Index $M$ of the highest eigenpair to compute.
            Maximum number of adaptive steps $n_{\mathrm{max}}$.
            Relative tolerance $\mathrm{tol}$.
            Length of the pause $\ell_{\mathrm{pause}}$.
      Output
            Approximated eigenpairs $(\lambda_{j,h},\phi_{j,h})$, for all $j=1,\dots,M$.
$n:=1$
2:$\mathrm{ct}_{\mathrm{pause}}:=\ell_{\mathrm{pause}}$
3:while $n<n_{\mathrm{max}}$ do
4:     Compute an approximation $u_{h}$ of the solution of the landscape function problem.
5:     Compute the error estimator $\eta_{\mathrm{land}}$.
6:     if $\mathrm{ct}_{\mathrm{pause}}=\ell_{\mathrm{pause}}$ then
7:         Compute an approximation of $(\lambda_{j,h},\phi_{j,h})$ computing the first $M$ eigenpairs at the bottom of the spectrum for the problem.
8:         Compute the error estimator $\eta_{\mathrm{eig},j}$, for all $j=1,\dots,M$.
9:         if $\max_{j\leq M}(\eta_{\mathrm{eig},j}^{2})<\mathrm{tol}$ then
10:              Return $(\lambda_{j,h},\phi_{j,h})$, for all $j=1,\dots,M$, and and exit.
11:         end if
12:         $\mathrm{ct}_{\mathrm{pause}}:=0$
13:     else
14:         $\mathrm{ct}_{\mathrm{pause}}:=\mathrm{ct}_{\mathrm{pause}}+1$
15:     end if
16:     Adapt $S(\mathcal{T})$ using $\eta_{\mathrm{land}}$ and $u_{h}$ by calling AdaptLandscape.
17:     $n:=n+1$
18:end while
In Table 3, Algorithm 8 is used to solve the problem already solved in Example 4.3 using different values for the pause length. The simulation for $\ell_{\mathrm{pause}}=0$ is equivalent to using Algorithm 3 and it is used as a reference to measure the reduction in CPU time. Since Algorithm 8 can only stop during an iteration with the eigensolver active, the total number of meshes varies depending on the value of $\ell_{\mathrm{pause}}$. For the same reason, also the DOF number on the final mesh varies with $\ell_{\mathrm{pause}}$. For any value of $\ell_{\mathrm{pause}}>0$ the reduction is great—up to 70%.
Pause Length N Meshes DOF CPU Time Reduction(%)
0 36 117805 30986.7 0
1 37 128366 19291.4 37.7428
2 37 128366 14276 53.9284
3 37 128366 11672.5 62.3306
4 36 117805 8940.47 71.1473
5 37 128366 9055.94 70.7747
Table 3. Comparison of performances for different values for the pause length.
Since the FE space in Algorithm 8 is adapted just using the landscape solution, it might seem a good idea to postpone any call to the eigensolver after a certain number of mesh refinements instead of pausing the eigensolver at intervals. Unfortunately, this approach does not reduce the total CPU time by much because the CPU time used by the eigensolver on the initial meshes is very moderate, see Figure 13.
6.2. Iterative solvers
In this section, we further reduce the CPU time by exploring different eigenvalue solvers. So far we used the ARPACK package [27] which is based on the Arnoldi method. ARPACK is very fast to compute multiple eigenpairs, however, in an adapting method, not all eigenpairs reach the desired tolerance on the same mesh. Therefore, it is common that on some meshes only a few of eigenpairs have not converged yet. ARPACK computes approximations of all eigenpairs, even of the ones already approximated well enough on previous meshes. In [20, 30], iterative eigenvalue solvers based on Picard’s and Newton’s methods are described. Even if these solvers might be slower than ARPACK in computing several eigenpairs, they can used to just compute a subset of the eigenpairs in the computed spectrum not necessarily contiguous in the spectrum.
In Figure 14, the number of not converged eigenpairs are reported for all computed meshes for the problem solved in Example 4.3. As can be seen, on the last few computed meshes, only a few eigenpairs have not converged yet but ARPACK is anyway computing the entire spectrum of 100 eigenpairs. These calls to the eigensolver are also the most expensive ones, as can be seen in Figure 13. Also in Figure 14, the CPU time for three different eigensolvers are tested from mesh 20 and onwards. ARPACK is still the fastest method until the last few meshes when the other methods are faster because only a small number of eigenpairs are computed. This suggests switching to an iterative solver when the number of not converged eigenpairs is small enough.
Figure 14. Left: Number of not converged eigenpairs on each adapted mesh. Right: CPU time comparison between different eigenvalue solvers
In Table 4, Algorithm 8 is used to solve the problem already solved in Example 4.3 using different values for the pause length and using the Picard eigensolver when the number of not converged eigenpairs is small than 30. The simulation for $\ell_{\mathrm{pause}}=0$ uses only ARPACK and it is equivalent to using Algorithm 3 and it is used as a reference to measure the reduction in CPU time. Compared to Table 3, we can see that the use of the iterative solver reduces further the CPU time.
Pause Length N Meshes DOF CPU Time Reduction(%)
0 (ARPACK only) 36 117805 30986.7 0
1 37 128366 19291.4 68.7414
2 37 128366 14276 77.0795
3 37 128366 11672.5 78.0287
4 36 117805 8940.47 80.9712
5 37 128366 9055.94 84.4589
Table 4. Comparison of performances for different values for the pause length and using an iterative solver on the last meshes.
7. Conclusions
We have introduced a new approach for adaptively approximating multiple eigenpairs of an elliptic operator $\mathcal{L}$ via finite element methods, where the adaptivity is driven not by the computed eigenpairs, but by the computed solution for a single source problem—the so-called landscape problem $\mathcal{L}u=1$. Although this approach was motivated by operators that induce strong spatial localization of many eigenvectors, and for which the landscape function has been seen to contain an impressive amount of information about these eigenvectors, we have demonstrated its utility on a broader range of problems. It compares favorably with more traditional approaches, both in terms of CPU time to reach a certain convergence criterion, and in terms of conceptual simplicity—refine based on approximation of a single function instead of figuring out how to incorporate information from a large collection of functions. We also presented a few modifications of the algorithm that reduce its computational cost even further. We believe that our landscape refinement approach provides an attractive alternative to traditional methods, particularly in cases where large numbers of eigenpairs are sought; and it can also be very effective in cases where a smaller cluster of eigenpairs is sought, as a sensible guide for refinement until the finite element space is “sufficiently rich”.
This work was partially supported by the National Science Foundation through NSF grant DMS-2012285 and NSF RTG grant DMS-2136228.